# A Generalized Approach to Online Convex Optimization

### v2

@@ 9 First order feedback to zeroth order feedback
While $\operatorname{STB}$ turns a semi-bandit algorithm into a bandit algorithm, $\operatorname{FOTZO}$ turns a general first order algorithm into a zeroth order algorithm. However, we note that even when ${\mathcal{A}}$ is semi-bandit, $\operatorname{FOTZO}({\mathcal{A}})$ is not a bandit algorithm. Instead it is a full-information zeroth order algorithm that uses a single query per iteration, but its query is not the same as its action. This is because $\operatorname{FOTZO}({\mathcal{A}})$ plays the same point ${\mathbf{x}}_{t}$ as the base algorithm ${\mathcal{A}}$, but then adds a noise before querying according to ${\mathcal{A}}^{\text{query}}$.
Remark 4.
We have shown in the previous sections that for online convex optimization, semi-bandit feedback is enough. However, we have included the $\operatorname{FOTZO}$ meta-algorithm, which converts a full-information algorithm to another full-information algorithm, since the results in this section apply also to non-convex online optimization problems.
Recall that OGD algorithm Zinkevich, (2003) is a deterministic algorithm designed for deterministic semi-bandit feedback. Hence we may apply Theorems 5 and 7, with $\alpha=\delta=T^{-1/4}$, to obtain the following result. This generalizes the result of Flaxman et al., (2005) to allow the feedback of the value oracle to be stochastic.
Corollary 3.
We have
${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\operatorname{STB}(\text{OGD})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(DB_{0}T^{3/4}\right).$
As mentioned in Remark 2, the Improved Ader algorithm of Zhang et al., 2018a is a deterministic algorithm designed for deterministic semi-bandit feedback. Hence we may apply Theorems 5 and 7, with $\alpha=\delta=T^{-1/4}$, to obtain the following result. This matches the SOTA result of Zhao et al., (2021) for dynamic regret in the bandit feedback setting and generalizes it to allow for stochastic feedback.
Corollary 4.
Let IA denote the algorithm mentioned above. Also let ${\mathbf{u}}\in{\mathcal{K}}^{T}$ and $P_{T}:={\mathbf{u}}\mapsto\sum_{t=1}^{T-1}\|{\mathbf{u}}_{t}-{\mathbf{u}}_{t+1}\|:{\mathcal{K}}^{T}\to{\mathbb{R}}$. Then we have
${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\operatorname{STB}(\text{IA})}({\mathbf{u}})[1,T]$ $=O\left(DB_{0}\left(1+\frac{P_{T}({\mathbf{u}})}{D}\right)^{1/2}T^{3/4}\right).$
Similarly, the SO-OGD algorithm Garber and Kretzu, (2022) is a deterministic algorithm designed for deterministic semi-bandit feedback with obtains a weakly adaptive regret bound of $O(DM_{1}T^{1/2})$ (Theorem 14 in Garber and Kretzu, (2022)) Hence we may apply Theorems 5 and 7, with $\alpha=\delta=T^{-1/4}$, to obtain the following result. This matches the result of Theorem 15 in Garber and Kretzu, (2022) and generalizes it to allow for stochastic feedback.
Corollary 5.
Let SO-OGD be the algorithm mentioned above. Then we have
$\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\operatorname{STB}(\text{SO-OGD})}({\mathcal{K}}_{*}^{T})[a,b]$ $=O\left(DB_{0}T^{3/4}\right).$
10 Acknowledgements
@@ Appendix B Proof of Theorem 2
Assume that $h:[0,1]\to{\mathbb{R}}$ is a high probability regret bound for $({\mathcal{A}},\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}))$. Using Inequality 1, we see that
@@ Appendix E Proof of Theorem 5
We also define ${\mathcal{B}}^{\prime}_{\theta}:=({\mathcal{B}}^{\prime}_{\theta,1},\cdots,{\mathcal{B}}^{\prime}_{\theta,T})$. Note that a specific choice of $\theta$ is necessary to make sure that the function returned by ${\mathcal{B}}^{\prime}_{\theta,t}$ is a deterministic function of ${\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{t}$ and not a random variable and therefore ${\mathcal{B}}^{\prime}_{\theta}$ belongs to $\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])$.
@@
${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}=\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}}\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1}),\theta\in\Omega^{\mathcal{Q}}}{\mathcal{R}}_{{\mathcal{B}}^{\prime}_{\theta}}^{{\mathcal{A}}}\leq\sup_{{\mathcal{B}}^{\prime}\in\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])}{\mathcal{R}}_{{\mathcal{B}}^{\prime}}^{{\mathcal{A}}}={\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])}^{{\mathcal{A}}}$
@@
$\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1}),\theta\in\Omega^{\mathcal{Q}}}{\mathcal{R}}_{{\mathcal{B}}^{\prime}_{\theta}}^{{\mathcal{A}}_{\omega}}\leq\sup_{{\mathcal{B}}^{\prime}\in\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])}{\mathcal{R}}_{{\mathcal{B}}^{\prime}}^{{\mathcal{A}}_{\omega}}={\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])}^{{\mathcal{A}}_{\omega}}$
@@
¯$\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])}^{{\mathcal{A}}}$
Appendix F Proof of Theorems 6 and 7
First we prove the following lemma which is a simple generalization of Observation 3 in Flaxman et al., (2005) to the construction of $\hat{{\mathcal{K}}}$ that we use here and is proved similarly.
@@ Appendix F Proof of Theorems 6 and 7
Otherwise, then let ${\mathbf{z}}={\mathbf{x}}+\alpha\frac{\Delta}{\|\Delta\|}$. According to Lemma 7 in Pedramfar et al., (2023), we have ${\mathbf{z}}\in{\mathcal{K}}$. We have ${\mathbf{y}}=\frac{\|\Delta\|}{\alpha}{\mathbf{z}}+\left(1-\frac{\|\Delta\|}{\alpha}\right){\mathbf{x}}$. Therefore, since $f$ is convex and bounded by $M_{0}$, we have
@@
where ${\mathbf{v}}$ is a random vector, taking its values uniformly from ${\mathbb{S}}^{1}\cap{\mathcal{L}}_{0}={\mathbb{S}}^{1}\cap(\operatorname{aff}({\mathcal{K}})-{\mathbf{z}})$, for any ${\mathbf{z}}\in{\mathcal{K}}$ and $k=\operatorname{dim}({\mathcal{L}}_{0})$. Since ${\mathcal{Q}}_{t}$ is a stochastic value oracle for $f_{t}$, according to Remark 4 in Pedramfar et al., (2023), $\hat{{\mathcal{Q}}}_{t}({\mathbf{x}})$ is an unbiased estimator of $\nabla\hat{f}_{t}({\mathbf{x}})$.  [^3] When using a spherical estimator, it was shown in Flaxman et al., (2005) that $\hat{f}$ is differentiable even when $f$ is not. When using a sliced spherical estimator as we do here, differentiability of $\hat{f}$ is not proved in Pedramfar et al., (2023). However, their proof is based on the proof for the spherical case and therefore the results carry forward to show that $\hat{f}$ is differentiable. Hence we have $\hat{B}\in\operatorname{Adv}_{1}^{\text{o}}(\hat{{\mathbf{F}}},\frac{k}{\delta}B_{0})$. Using Equation 3 and the definition of the Algorithm 3, we see that the responses of the queries are the same between the game $({\mathcal{A}},\hat{{\mathcal{B}}})$ and $({\mathcal{A}}^{\prime},{\mathcal{B}})$. It follows that the sequence of actions $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{T})$ in $({\mathcal{A}},\hat{{\mathcal{B}}})$ corresponds to the sequence of actions $({\mathbf{x}}_{1}+\delta{\mathbf{v}}_{1},\cdots,{\mathbf{x}}_{T}+\delta{\mathbf{v}}_{T})$ in $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
@@
To prove the result in the Lipschitz case, we note that, according to Lemma 3 in Pedramfar et al., (2023), we have $|\hat{f_{t}}({\mathbf{x}}_{t})-f_{t}({\mathbf{x}}_{t})|\leq\delta M_{1}$. By using Lipschitz property for the pair $({\mathbf{x}}_{t},{\mathbf{x}}_{t}+\delta{\mathbf{v}}_{t})$, we see that
@@
$\leq\delta M_{1}T+\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{u}}}_{t})$ (Lemma 3 in Pedramfar et al., (2023))
@@
The proof of the bounds for this case is similar to the previous case. As before, we see that the responses of the queries are the same between the game $({\mathcal{A}},\hat{{\mathcal{B}}})$ and $({\mathcal{A}}^{\prime},{\mathcal{B}})$. It follows from the description of Algorithm 2 that the sequence of actions $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{T})$ in $({\mathcal{A}},\hat{{\mathcal{B}}})$ corresponds to the same sequence of actions in $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
@@
To obtain the same bound as before, we note that in the convex case, instead of Inequality 5, we have
@@
and, in the Lipschitz case, instead of Inequality F, we have

### v3

@@ 11 Applications
A careful review of Algorithms 6 and 7 in (?), together with their Lemmas 12 and 13 and Theorems 14 and 15 reveals that they added an extra layer of complexity to their base algorithm in order to be able to move from semi-bandit to bandit feedback. Specifically, if we instead use our Theorem 7, we can just drop $\delta^{\prime}$ from Algorithms 6 and 7 and Lemmas 12 and 13 (by setting it to zero) to obtain simpler and clearer algorithms, statements and proofs. Then their Theorem 15 will be an immediate corollary of applying our Theorem 7 to their previous results and the 3 pages of proof in their Appendix I will not be needed.
12 Conclusions
This paper presents a comprehensive framework for the analysis of meta-algorithms in online learning, offering a simplified approach to understanding and extending existing results in the field. By demonstrating the equivalence between online linear and convex optimization algorithms, particularly in the context of semi-bandit feedback, we provide a versatile tool for addressing a wide range of optimization problems. Moreover, our findings pave the way for the development of novel meta-algorithms, enabling the seamless conversion of algorithms across different settings. In addition to recovering many results in the literature with simplified proofs, we show that adaptive regret for convex function is $O(\sqrt{T})$ and static regret for strongly convex function is $O(\log T)$ with deterministic zeroth order feedback.
This work opens various research opportunities to delve into the interplay between online linear and convex optimization algorithms. While it streamlines results across numerous scenarios, we anticipate that this approach will yield novel insights into some unresolved problems moving forward.
13 Acknowledgments
@@ B Proof of Theorem 2
Assume that $h:[0,1]\to{\mathbb{R}}$ is a high probability regret bound for $({\mathcal{A}},\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}))$. Using Inequality 4, we see that
@@ E Proof of Theorem 5
We also define ${\mathcal{B}}^{\prime}_{\theta}:=({\mathcal{B}}^{\prime}_{\theta,1},\cdots,{\mathcal{B}}^{\prime}_{\theta,T})$. Note that a specific choice of $\theta$ is necessary to make sure that the function returned by ${\mathcal{B}}^{\prime}_{\theta,t}$ is a deterministic function of ${\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{t}$ and not a random variable and therefore ${\mathcal{B}}^{\prime}_{\theta}$ belongs to $\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])$.
@@
${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}=\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}}\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1}),\theta\in\Omega^{\mathcal{Q}}}{\mathcal{R}}_{{\mathcal{B}}^{\prime}_{\theta}}^{{\mathcal{A}}}\leq\sup_{{\mathcal{B}}^{\prime}\in\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}{\mathcal{R}}_{{\mathcal{B}}^{\prime}}^{{\mathcal{A}}}={\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}^{{\mathcal{A}}}$
@@
$\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1}),\theta\in\Omega^{\mathcal{Q}}}{\mathcal{R}}_{{\mathcal{B}}^{\prime}_{\theta}}^{{\mathcal{A}}_{\omega}}\leq\sup_{{\mathcal{B}}^{\prime}\in\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}{\mathcal{R}}_{{\mathcal{B}}^{\prime}}^{{\mathcal{A}}_{\omega}}={\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}^{{\mathcal{A}}_{\omega}}$
@@
¯$\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}^{{\mathcal{A}}}$
F Proof of Theorems 6 and 7
First we prove the following lemma which is a simple generalization of Observation 3 in (?) to the construction of $\hat{{\mathcal{K}}}$ that we use here and is proved similarly.
@@ F Proof of Theorems 6 and 7
Otherwise, then let ${\mathbf{z}}={\mathbf{x}}+\alpha\frac{\Delta}{\|\Delta\|}$. According to Lemma 7 in (?), we have ${\mathbf{z}}\in{\mathcal{K}}$. We have ${\mathbf{y}}=\frac{\|\Delta\|}{\alpha}{\mathbf{z}}+\left(1-\frac{\|\Delta\|}{\alpha}\right){\mathbf{x}}$. Therefore, since $f$ is convex and bounded by $M_{0}$, we have
@@
where ${\mathbf{v}}$ is a random vector, taking its values uniformly from ${\mathbb{S}}^{1}\cap{\mathcal{L}}_{0}={\mathbb{S}}^{1}\cap(\operatorname{aff}({\mathcal{K}})-{\mathbf{z}})$, for any ${\mathbf{z}}\in{\mathcal{K}}$ and $k=\operatorname{dim}({\mathcal{L}}_{0})$. Since ${\mathcal{Q}}_{t}$ is a stochastic value oracle for $f_{t}$, according to Remark 4 in (?), $\hat{{\mathcal{Q}}}_{t}({\mathbf{x}})$ is an unbiased estimator of $\nabla\hat{f}_{t}({\mathbf{x}})$.  [^3] When using a spherical estimator, it was shown in (?) that $\hat{f}$ is differentiable even when $f$ is not. When using a sliced spherical estimator as we do here, differentiability of $\hat{f}$ is not proved in (?). However, their proof is based on the proof for the spherical case and therefore the results carry forward to show that $\hat{f}$ is differentiable. Hence we have $\hat{B}\in\operatorname{Adv}_{1}^{\text{o}}(\hat{{\mathbf{F}}},\frac{k}{\delta}B_{0})$. Using Equation 6 and the definition of the Meta-algorithm 3, we see that the responses of the queries are the same between the game $({\mathcal{A}},\hat{{\mathcal{B}}})$ and $({\mathcal{A}}^{\prime},{\mathcal{B}})$. It follows that the sequence of actions $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{T})$ in $({\mathcal{A}},\hat{{\mathcal{B}}})$ corresponds to the sequence of actions $({\mathbf{x}}_{1}+\delta{\mathbf{v}}_{1},\cdots,{\mathbf{x}}_{T}+\delta{\mathbf{v}}_{T})$ in $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
@@
To prove the result in the Lipschitz case, we note that, according to Lemma 3 in (?), we have $|\hat{f_{t}}({\mathbf{x}}_{t})-f_{t}({\mathbf{x}}_{t})|\leq\delta M_{1}$. By using Lipschitz property for the pair $({\mathbf{x}}_{t},{\mathbf{x}}_{t}+\delta{\mathbf{v}}_{t})$, we see that
@@
$\leq\delta M_{1}T+\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{u}}}_{t})$ (Lemma 3 in (?))
@@
The proof of the bounds for this case is similar to the previous case. As before, we see that the responses of the queries are the same between the game $({\mathcal{A}},\hat{{\mathcal{B}}})$ and $({\mathcal{A}}^{\prime},{\mathcal{B}})$. It follows from the description of Meta-algorithm 2 that the sequence of actions $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{T})$ in $({\mathcal{A}},\hat{{\mathcal{B}}})$ corresponds to the same sequence of actions in $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
@@
To obtain the same bound as before, we note that in the convex case, instead of Inequality 8, we have
@@
and, in the Lipschitz case, instead of Inequality 9, we have
@@
G Proof of Theorem 8
This proof is similar to the proofs of Theorems 6 and 7.
Proof.
Note that any realized adversary ${\mathcal{B}}\in\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})$ may be represented as a sequence of functions $(f_{1},\cdots,f_{T})$. For such realized adversary ${\mathcal{B}}$, we define $\hat{B}$ to be the realized adversary corresponding to $(\hat{f}_{1},\cdots,\hat{f}_{T})$ with the stochastic gradient oracles
$\hat{{\mathcal{Q}}}_{t}({\mathbf{x}}):=\frac{k}{2\delta}\left(f_{t}({\mathbf{x}}+\delta{\mathbf{v}})-f_{t}({\mathbf{x}}-\delta{\mathbf{v}})\right){\mathbf{v}},$ (12)
where ${\mathbf{v}}$ is a random vector, taking its values uniformly from ${\mathbb{S}}^{1}\cap{\mathcal{L}}_{0}={\mathbb{S}}^{1}\cap(\operatorname{aff}({\mathcal{K}})-{\mathbf{z}})$, for any ${\mathbf{z}}\in{\mathcal{K}}$ and $k=\operatorname{dim}({\mathcal{L}}_{0})$. Since ${\mathcal{Q}}_{t}$ is a stochastic value oracle for $f_{t}$, according to Lemma 5 in (?), $\hat{{\mathcal{Q}}}_{t}({\mathbf{x}})$ is an unbiased estimator of $\nabla\hat{f}_{t}({\mathbf{x}})$. Moreover, we have
$\|\frac{k}{2\delta}\left(f_{t}({\mathbf{x}}+\delta{\mathbf{v}})-f_{t}({\mathbf{x}}-\delta{\mathbf{v}})\right){\mathbf{v}}\|\leq\frac{k}{2\delta}M_{1}\|({\mathbf{x}}+\delta{\mathbf{v}})-({\mathbf{x}}-\delta{\mathbf{v}})\|\leq kM_{1}.$
Hence we have $\hat{B}\in\operatorname{Adv}_{1}^{\text{o}}(\hat{{\mathbf{F}}},kM_{1})$. Using Equation 12 and the definition of the Meta-algorithm 4, we see that the responses of the queries are the same between the game $({\mathcal{A}},\hat{{\mathcal{B}}})$ and $({\mathcal{A}}^{\prime},{\mathcal{B}})$. It follows that the sequence of actions $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{T})$ in $({\mathcal{A}},\hat{{\mathcal{B}}})$ corresponds to the same sequence of actions in $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
According to Lemma 3 in (?), we have $|\hat{f_{t}}({\mathbf{x}}_{t})-f_{t}({\mathbf{x}}_{t})|\leq\delta M_{1}$. On the other hand, we have
$\leq\delta M_{1}T+\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{u}}}_{t})$ (Lemma 3 in (?))
$=\delta M_{1}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}\left(\left(1-\frac{\delta}{r}\right){\mathbf{u}}_{t}+\frac{\delta}{r}{\mathbf{c}}\right)$ (Definition of $\hat{{\mathcal{U}}}$)
$=\delta M_{1}T+\min_{{\mathbf{u}}\in{\mathcal{K}}}\sum_{t=a}^{b}f_{t}\left({\mathbf{u}}_{t}+\frac{\delta}{r}({\mathbf{c}}-{\mathbf{x}})\right)$
$\leq\delta M_{1}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(f_{t}({\mathbf{u}}_{t})+\frac{2\delta M_{1}D}{r}\right)$ (Lipschitz)
$=\left(1+\frac{2D}{r}\right)\delta M_{1}T+\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})$
Therefore, using Equation 13, we see that
${\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}^{\prime}}-{\mathcal{R}}_{\hat{{\mathcal{B}}}}^{{\mathcal{A}}}$ $\leq\delta M_{1}T+\left(1+\frac{2D}{r}\right)\delta M_{1}T=\left(2+\frac{2D}{r}\right)\delta M_{1}T.$
Therefore, we have
${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}$ $=\sup_{{\mathcal{B}}\in\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}^{\prime}}$
$\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}{\mathcal{R}}_{\hat{{\mathcal{B}}}}^{{\mathcal{A}}}+\left(2+\frac{2D}{r}\right)\delta M_{1}T$
$\leq\sup_{{\mathcal{B}}\in\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}{\mathcal{R}}_{\hat{{\mathcal{B}}}}^{{\mathcal{A}}}+\left(2+\frac{2D}{r}\right)\delta M_{1}T$
$\leq{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}(\hat{{\mathbf{F}}},kM_{1})}^{{\mathcal{A}}}+\left(2+\frac{2D}{r}\right)\delta M_{1}T.\qed$